OpenStudy (anonymous):

Consider the function f(x) = 8 − 3x2. Find the value of f(x) when x = 3 and select the correct answer below. –19 –10 26 35

OpenStudy (anonymous):

This one is going to be hard for me to explain cause i always show my work

OpenStudy (whpalmer4):

Is that \[f(x) = 8 -3 x^2\] ?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Okay, so what does \(f(3) = \) just replace \(x\) in the equation with \(3\)

OpenStudy (whpalmer4):

\[f(3) = 8-3(3)^2 = \]

OpenStudy (anonymous):

-19

OpenStudy (whpalmer4):

Yep!